Show that the function f :R -> R defined by f(x) = 2x+ 7 has an inverse by applying the inverse function theorem. Find its inverse also 


1
Expert's answer
2021-05-11T07:01:12-0400

We have given the function,


"f: R \rightarrow R" defined by "f(x) = 2x+7"


Since, the given function is bijective because for every y there exist a unique x


"x = \dfrac{y-7}{7}" , such that "f(x) = y"


In general we can say that "R \rightarrow R"


"f(x) = ax+b, a \ne 0"


Hence, we can say that inverse exist for the given function "f(x)".


Calculation of "f^{-1}(x)."


"f(x)= 2x+7"


"y = 2x+7"


"x = \dfrac{y-7}{2}"


where, "x = f^{-1}(x)."


Hence, "f^{-1}(x) = \dfrac{y-7}{2}."